OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (2, -5) with a slope of -3. y = -3x - 13 y = -3x + 11 y = -3x + 13 y = -3x + 1

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

The general slope intercept form is y = mx+b the m is the slope (think "m" for mountain, mountains have slopes) since the slope is -3, we know that m = -3 so the equation is in the form y = -3x + b

jimthompson5910 (jim_thompson5910):

the line passes through (2,-5) (x,y) = (2,-5) x = 2 y = -5 y = -3x + b y = -3*2 + b ... replace x with 2 y = -6 + b -5 = -6 + b ... replace y with -5 now solve for b

OpenStudy (anonymous):

now how would I do that, I don't quite get how to solve for b.

jimthompson5910 (jim_thompson5910):

think of -6+b as b-6

jimthompson5910 (jim_thompson5910):

-5 = b-6 try to move that -6 over to isolate b

OpenStudy (usukidoll):

have b by itself.. what do you need to do to make this happen?

OpenStudy (anonymous):

subtract -6 from the sides?

jimthompson5910 (jim_thompson5910):

close

jimthompson5910 (jim_thompson5910):

oh nvm, I was thinking of something else yeah you are correct

jimthompson5910 (jim_thompson5910):

subtracting -6 from both sides is the same as adding 6 to both sides

jimthompson5910 (jim_thompson5910):



jimthompson5910 (jim_thompson5910):